The price elasticity for rice is estimated to be 0.4 and the income elasticity is 0.8. At a price of $0.40 per pound and a per capita income of $20,000, the demand for rice is 50 million tons per year.
Is rice an inferior good, a necessity, or a luxury? Explain.
If per capita income increases to $20,500, approximately what will be the quantity demanded rice?
1
Expert's answer
2020-11-23T10:21:47-0500

Rice is a basic food for billion people, so it is a necessity good

"dI=20,500-20,000=500"

"dQ=500*0.8=400\ tons"

New Q: "50,000,000+400=50,000,400\ tons"